Show how you would solve by Substitution:
2x + y = 0
x – 3y = -7

Show how you would solve by Elimination: 3x + 4y = 1
5x – 3y = 21

Answer by checkley77(12844)

2x + y = 0 MULTIPLY BY 3 & ADD
x – 3y = -7
-6X+3Y=0
-------------------
-5Y=-7
Y=-7/-5
Y=7/5 ANS.
2X+7/5=0
2X=-7/5
X=-7/5/2
X=-7/5*1/2
X=-7/10 ANS.
PROOF:
2(-7/10)+7/5=0
-14/10+7/5=0
-7/5+7/5=0
0=0